OpenStudy (anonymous):

A woman driving a 2.0 x 10^3 kg car at 15m/s fully applies the brakes 50.m from a stop light. If the car stops 5.0m before the light, what is the magnitude of the average force applied by the brakes? (4.5 x 10^3)??

OpenStudy (stormfire1):

\[Vi=15 m/s\]\[Vf = 0m/s\]\[Xf-Xi=0m-45m=-45m\]Use the following kinematic equation to find the (negative) acceleration caused by her braking: \[V_f^2-V_i^2=2a(X_f-X_i)\]Solve for a to get:\[a=\frac{(-15m/s)^2}{-90m}=-2.5 m/s^2\]Now find the force:\[F=ma=(2000kg)(-2.5 m/s^2)=-5000N\] The magnitude of the force applied by her brakes is therefore 5000N.